The larger of two supplementary angles is 40 degrees less than 3 times the smaller. Find both angles

Answer by checkley71(8403)

X+3X-40=180
4X=180+40
4X=220
X=220/4
X=55 DEGREES FOR THE SMALLER ANGLE.
180-55=125 FOR THE LARGER ANGLE.
55+3*55-40=180
55+165-40=180
220-40=180
180=180